OpenStudy (anonymous):

factor the polynomial using a special factorization: 16x^3-2y^3

jhonyy9 (jhonyy9):

you need using a*3 - b*3 =(a-b)(a*2 +ab +b*2)

jhonyy9 (jhonyy9):

after this form 2(8x*3 -y*3) = 2(2x -y)(4x*2 +2xy +y*2)